OpenStudy (anonymous):

The Pattern is that (Leg 1) ^2= Leg 2+ Hyp. This only works if Leg 1 is an odd integer. Prove that this is true for all odd integers.

OpenStudy (anonymous):

i dont understand the question =/ yo myininaya :)

myininaya (myininaya):

if Leg 1=2k+1, then (Leg 1)^2=4k^2+4k+1 so we have 4k^2+4k+1=leg 2 +hyp now we got to see if this is possible when it says prove for all odd integers does that mean leg 2 and hyp are odd integers? .... hey joe

OpenStudy (anonymous):

no just leg 1 is an odd integer

OpenStudy (anonymous):

one of either leg 2 or the hyp would need to be even, the other odd.

OpenStudy (anonymous):

yes...

myininaya (myininaya):

(3,4,5) 9=4+5 so leg 1 is also have to be the smallest leg i believe

myininaya (myininaya):

because 16 doesn't equal 3+5

myininaya (myininaya):

oh and 4 is not odd

OpenStudy (anonymous):

but hows does this prove tru for all odd integers?

myininaya (myininaya):

it doesn't i was just thinking

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

this isnt a true statement =/ check out this pythagorean triple \[161^2+240^2=289^2\] \[161^2=25921\]

OpenStudy (anonymous):

You are looking at pythagorean triples i assume (since it was really never said). So they must satisfy the equation: \[(leg_1)^2+(leg_2)^2=(hyp)^2\iff(leg)^2=(hyp)^2-(leg_2)^2\]\[\iff (leg_1)^2=(hyp+leg_2)(hyp-leg_2)\] So, in order for your statement to be true, it must follow that:\[hyp-leg_2=1\] ,which is the case for the "smaller" pythagorean triples, but not the larger ones.